OpenStudy (anonymous):

Graph this circle..or at least gimme some coordinates: x^2+y^2-3x+8y=28

OpenStudy (mimi_x3):

Complete the square. x^2 - 3x + y^2 + 8y - 28 = 0

OpenStudy (anonymous):

wad do u do then?

OpenStudy (mimi_x3):

Complete the square first.

OpenStudy (saifoo.khan):

Help?

OpenStudy (anonymous):

yes please

OpenStudy (saifoo.khan):

im going to delete it all. xD

OpenStudy (anonymous):

if u do ima tell a mod

OpenStudy (mimi_x3):

Im not going to do it for you

OpenStudy (mimi_x3):

In fact, saif can help

OpenStudy (saifoo.khan):

i dont help bad guys.

OpenStudy (anonymous):

okay..if i complete the square...what happens next? i have to have it in (x-h)^2+(y+k)^2=r^2

OpenStudy (mimi_x3):

You should help, i remember giving you information on how to do it

OpenStudy (anonymous):

pretty pretty please with a cherry on top

OpenStudy (mimi_x3):

Complete the square and you will get it into that form.

OpenStudy (anonymous):

how do u complete the square with that?

OpenStudy (saifoo.khan):

we can't.

OpenStudy (mimi_x3):

first x^3 - 3x ( x + b/2)^2 - (b/2)^2 Goodman, try it

OpenStudy (anonymous):

i dont understand..we learned it a different way...

OpenStudy (saifoo.khan):

Awesome.

OpenStudy (mimi_x3):

What's the other way?

OpenStudy (anonymous):

well...we usually dont "complete the square" in this sitution...im conpoosed

OpenStudy (anonymous):

*confused

OpenStudy (mimi_x3):

well, what do you do? i think that there's another method but i forgot

OpenStudy (anonymous):

i hate circles...anyways, we only used completing the square wen dealing with un-factorable questions

OpenStudy (mimi_x3):

Completeing the square is possible. Saif, can help

OpenStudy (saifoo.khan):

i can help good guys

OpenStudy (anonymous):

fine..i am good@saifoo <= i called u ur real name

OpenStudy (anonymous):

now help me please :P

OpenStudy (saifoo.khan):

http://www.youtube.com/watch?v=9TvAAuksFA4

OpenStudy (anonymous):

i cant use youtube..iam on my computer, not laptop..cuz our wifi has been cut off due to my brother breaking our router..can u explain it here?

OpenStudy (mimi_x3):

(x + 3/2)^2 - (3/2)^2 + (y+4)^2 - (4)^2 - 28 = 0 Simplify and collect the like terms.

OpenStudy (saifoo.khan):

is it a crime to use youtube from your pc?

OpenStudy (mimi_x3):

w8, ignore simplify

OpenStudy (anonymous):

its ram and harddrive is filled to its limits..it is currently very hard to use OS without getting stuck now and then...so wen i use youtube, its like a death sentence to my internet..

OpenStudy (anonymous):

SO HOW DO I SOLVE THIS???

OpenStudy (mimi_x3):

I JUST SHOWED YOU !! please don't be lazy like saifi ALWAYS getting the answers

OpenStudy (saifoo.khan):

Answers Rules!

OpenStudy (anonymous):

im perty busy rite now..i gotta finish this then i have to read a stupid book..and ya..ANSWERS DO RULE..

OpenStudy (anonymous):

i dont mind...well not now..can u jus gimmie the answers?

OpenStudy (mimi_x3):

Fine, use wolfram

OpenStudy (anonymous):

that site doesnt help me..i use it all the time..

OpenStudy (saifoo.khan):

good people get help, not bad.

OpenStudy (mimi_x3):

Answers given :D Just be lazy like saif.

OpenStudy (anonymous):

i will be rite bak...ughhh

OpenStudy (anonymous):

help me now..im back..and i needa go soon..fast and quick

OpenStudy (anonymous):

u know wat..i dont care no more...this is getting stupid, im jus gonna copy wolfram, thanks for the help anyways

OpenStudy (mimi_x3):

LOL, well, i did help, you just didnt do it

OpenStudy (anonymous):

yes mimi..i appreciate ur help a lot!! thank you..i think its jus my head that doesnt wanna do nothing but sit and be lazy...oh well..thank u anyways oh and thanks seafood, whoops..saifoo, for giving me some sort of weird inspiration for jus copying..i think i shood be thankful..lols